OpenStudy (anonymous):

I have the answer can someone just see if I'm right. f(n)=3n^2-2 and g(n)= n+1 The domain of f(n)/g(n)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I am not sure if I did it right do you have an answer

OpenStudy (anonymous):

you are trying to find the domain of a rational function right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

from what i can see your domain is every value apart from -1

OpenStudy (anonymous):

because -1+1=0, that will lead to the 1/0= undefined at ur level

OpenStudy (anonymous):

so -1 does not satisfy this equation

OpenStudy (anonymous):

So is the answer then just 3n^2-2

OpenStudy (anonymous):

no dear

OpenStudy (anonymous):

this is not about simplifying the expression

OpenStudy (anonymous):

this is about what numbers can i plug in n

OpenStudy (anonymous):

yes it is Find and simplify

OpenStudy (anonymous):

Question is ((3n^2-2)/(n+1))

OpenStudy (anonymous):

yes from what i am reading its says find the domain

OpenStudy (anonymous):

That is true I am really confused right now

OpenStudy (anonymous):

\[\frac{ 2n^2 -2}{ n+1}\]

OpenStudy (anonymous):

don't be

OpenStudy (anonymous):

Put a 3 instead of a 2 for 3n^2

OpenStudy (anonymous):

\[\frac{ 3n^2 -2}{ n+1}\]

OpenStudy (anonymous):

yes sorry it would have been easier if it was 2 anyways

OpenStudy (anonymous):

Yes that is the equation now find and simplify also The doamin of f(n)/g(n)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Do you have an answer so I can see if it matches mine

OpenStudy (anonymous):

i gave u the answer for the domain

OpenStudy (anonymous):

now u added a new part so i will solve it for u

OpenStudy (anonymous):

What -1

OpenStudy (anonymous):

recall x-axis right

OpenStudy (anonymous):

yese

OpenStudy (anonymous):